A cooling fan is turned off when it is running at 800 rev/min . It turns 700 revolutions before it comes to a stop.

Part A: What was the fan's angular acceleration, assumed constant?

Part B: How long did it take the fan to come to a complete stop?
Answer #1

Ans=

The angular acceleration will be calculated by this equation =  ω2f - ω2i = 2αΔθ

= 0 - (83.73 rad/s)2 = 2*α*4396 rad

  α = - 0.79 rad/s2

Time taken to complete stopping of the fan = t = (ωf − ωi) / α

t = ( 0 - 83.73) / 0.79 = 106 sec